calculate the percent recovery if 7.7g of fat are collected from 4.1g of nutmeg​. nutmeg contains approximately 36% fat by weight.
Answer #1

Mass of pure compound within the impure sample = (36 / 100) * 4.1 g = 1.476 g

Percent recovery = Mass of pure compound within the impure sample * 100

                                     Total mass of the impure sample     

                         = (1.476 / 7.7) *100

Percent recovery = 19.16 %.